# Distributed Split Computing Using Diffusive Metrics for UAV Swarms

### v1

@@ V-C Effect of Network Distribution
Figure 5: Performance metrics under varying mission area sizes. Panel (a) shows average latency, (b) depicts remaining GFLOPs, and (c) presents the overall figure of merit.
V-D Early-Exit Mechanism
@@
Figure 6: Early-exit mechanism performance as a function of the number of workers. Panel (a) shows average accuracy, (b) depicts average latency, (c) illustrates remaining GFLOPs, (d) shows the fairness index, (e) presents energy per task, and (f) shows the overall figure of merit.
VI Conclusion
@@
The proposed approach opens several avenues for future work. One potential extension involves incorporating advanced mobility models and real flight data to capture even more realistic UAV trajectories. Additionally, while we primarily focused on split learning at the inference level, further exploration could integrate this method with federated training, enabling on-the-fly updates to model parameters in addition to distributing intermediate layers. Another intriguing direction involves dynamically adjusting the structure of the neural network layers themselves—employing model compression or pruning techniques—to further reduce communication overhead when bandwidth constraints become severe. We expect that, taken together, these refinements will push the frontier of distributed intelligence in swarms, enabling UAV networks of unprecedented scale and complexity to function reliably in even the most challenging environments.
@@ Acknowledgment
This work is supported by The Scientific and Technological Research Council of Turkey (TUBITAK) 1515 Frontier R&D Laboratories Support Program for BTS Advanced AI Hub: BTS Autonomous Networks and Data Innovation Lab. Project 5239903 and Istanbul Technical University, Department of Scientific Research Projects (ITU-BAP, 45375), and was partially supported by the PRIN 2022 project RAIN4C: “Reliable Aerial and satellIte Networks: joint Communication, Computation, Caching for Critical scenarios” (Id: 20227N3SPN, CUP: J53D23007020001).

### v2

@@ 5.3 Effect of Network Distribution
Figure 6: Early-exit mechanism performance as a function of the number of workers. Figure (a) shows average accuracy, (b) depicts average latency, (c) illustrates remaining GFLOPs, (d) shows the fairness index, (e) presents energy per task, and (f) shows the overall figure of merit.
5.4 Early-Exit Mechanism
@@
6 Conclusion
@@
The proposed approach opens several avenues for future work. One potential extension involves incorporating advanced mobility models and real flight data to capture even more realistic UAV trajectories. Additionally, while we primarily focused on split learning at the inference level, further exploration could integrate this method with federated training, enabling on-the-fly updates to model parameters in addition to distributing intermediate layers. Another intriguing direction involves dynamically adjusting the structure of the neural network layers themselves, employing model compression or pruning techniques, to further reduce communication overhead when bandwidth constraints become severe. We expect that, taken together, these refinements will push the frontier of distributed intelligence in swarms, enabling UAV networks of unprecedented scale and complexity to function reliably in even the most challenging environments.
@@ Acknowledgment
This work is supported by Istanbul Technical University, Department of Scientific Research Projects (ITU-BAP, 45375), and was partially supported by the PRIN 2022 project RAIN4C: “Reliable Aerial and satellIte Networks: joint Communication, Computation, Caching for Critical scenarios” (Id: 20227N3SPN, CUP: J53D23007020001).